OpenStudy (anonymous):

Find an othogonal matrix P such that (P-transpose*A*P) diagonalizes matrix A. I use the usual eigenvalue-vector method but the eigenvecoters i get are dodge. could someone please go through how they get this!¡ Thanks in advance!!

OpenStudy (anonymous):

A=(0 3 0 3 0 4 0 4 0) this might help -.- hahah

OpenStudy (phi):

did you find the eigenvalues? find the characteristic equation \[\left| \left[\begin{matrix}-\lambda & 3 & 0 \\ 3 & -\lambda & 4 \\ 0& 4 & -\lambda \end{matrix}\right]\right|=0\] you get \[−λ(λ^2 -16) -3(−3λ) = 0 \\ λ( -λ^2 +16+9)=0 \\ λ( -λ^2 +25)=0 \] solve for lambda. what do you get ?

OpenStudy (anonymous):

got plus and minus 5, then i did the whole eigenvector thing but thats where i com across the problem. Plug in 5 and X1, X2, X3 are all equal zero, same thing for -5 and 0 isn't an option... Probably solving the equations incorrectly!

OpenStudy (anonymous):

doing the rref i get a value in each pivot position with 0's in the rest, do i take those values as the eigenvectors? go through it for me cause I'm a tad at a loss! thanks

OpenStudy (phi):

yes, the eigenvalues are 5, -5, and 0 notice the equation that defines the eigenvalue/vector is \[ A x = \lambda x\\ A x - \lambda x=0 \\ \left(A -λ I\right) x = 0 \] which tells you that the eigenvector is in the null space of the matrix (A - λI) to find the eigenvectors, find the null space of this matrix. for each eigenvalue, create the matrix \[ \left[\begin{matrix}-\lambda & 3 & 0 \\ 3 & -\lambda & 4 \\ 0& 4 & -\lambda \end{matrix} \right]\] and use Gauss-Jordan elimination https://en.wikipedia.org/wiki/Kernel_(matrix)

OpenStudy (anonymous):

Yep, I do this part and get two vectors (each) for -5 and 5, 0 is not an option. I pick the three which are L.I. and come out with this B=(3 0 0 5 -5 4 0 4 5) Could you check if this is correct? Once I've got that I have to orthonormalize B in order to diagonlize A right?

OpenStudy (phi):

You should get 3 eigenvectors, including an eigenvector for eigenvalue 0 for example, if you start with 0 3 0 3 0 4 0 4 0 which is for lambda=0. Perform gauss-jordan elimination. we can swap rows 3 0 4 0 3 0 0 4 0 divide the first and second rows by 3 1 0 4/3 0 1 0 0 4 0 add -4 * 2nd row to the 3rd row 1 0 4/3 0 1 0 0 0 0 if this is multiplies the vector x y z, with z=1. you find y=0, x= -4/3 the eigenvector is [-4/3 0 1] multiply by 3 to get v= [- 4 0 3]^T as a test multiply A v to get [0 0 0] so Av = 0 v works.

OpenStudy (phi):

notice that matrix A is symmetric. that means its eigenvectors are orthogonal. but we do need to make them unit length. v = v= [-4 0 3]/sqrt(25)= [-4/5 0 3/5] you must do the same thing for eigenvalues 5 and -5 the corresponding eigenvectors will be the columns of matrix P and P^T A P = a diagonal matrix

OpenStudy (anonymous):

Awesome dude! been going round in circles for a couple days. thanks¡¡¡!!!